OpenStudy (anonymous):

For the following polynomial functions give their degree, leading coefficient, and constant term: 1. f(x)= -12x^5 -4x^4 +3x^3- 10x=6 pls help

OpenStudy (anonymous):

degree is 5(The highest power of x in a polynomial). Leading coeffecient is the number accompanying the highest power of x. Here it is -12. Constant is the term without any x which is -6(Taking 6 to the other side).

OpenStudy (anonymous):

thank u :) could u explain in an easier way so i may do others by myself :)

OpenStudy (anonymous):

wait lol i understand buddy thank u :) my b aint really read it all :) thanks

OpenStudy (anonymous):

Just see the powers of x. Here you have like the powers of x as 5,4,3,1,0. So the highest power is the degree of the polynomial which is 5. So now see x^5 term. The number accompanying it is the leading coeffecient .Here you have -12x^5. So it is -12.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

wiat the problem isnt =6 its + 6 my bad so does that mean its +6 for constant :) ??

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

ok i got another its differnet. F(x)= x(x-9)^2 its says to find all zeroes and describe mulitplicity

OpenStudy (anonymous):

The zeroes mean the roots of the equation F(x)=0. Now it is clear that at x=0 and x=9, F(x)=0.

OpenStudy (anonymous):

how did u get that? i have no idea how u got that

OpenStudy (anonymous):

If you have an equation like (x-a)(x-b)(x-c)(x-d)=0. Then all possible roots are x=a,x=b,x=c,x=d.(if you put x=a a-a=0. and the entire thing reduces to 0.) . Similarly in yours (x-0)(x-9)(x-9)=0. So the roots are 0,9.

OpenStudy (anonymous):

ok that makes sense so how would i start on x^2(x+2)^2(x^2-1)

OpenStudy (anonymous):

again (x-0)(x-0)(x+2)(x+2)(x^2-1)=0 That means x=0,x=-2 or x^2-1=0 or x^2=1which means x=1,x=-1/ So the solutions are x=-2,-1,0,1

OpenStudy (anonymous):

uhmmm ok i think